OpenStudy (anonymous):

The graph of f(x)=2^x+4 shifts six units o the right when it is replaced with the graph of f(x)=2^x-k. What is the value of k? a. 2 b. 4 c. 6 d. 10

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

what number is 6 units from 4 in the negative direction...?

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

thats it.. so since the expression is x - k the value is k = 2 x + 4 - 6 = x - 2 so k = 2

OpenStudy (anonymous):

Thank you!